### v1

# Learning Japanese with Jouzu: Interaction Outcomes with Stylized Dialogue Fictional Agents

Abstract.
This study investigates how stylized, voiced agents shape user interaction in a multimodal language learning environment. We conducted a mixed-methods evaluation of 54 participants interacting with anime-inspired characters powered by large language models and expressive text-to-speech synthesis. These agents responded in Japanese character language, offering users asynchronous, semi-structured conversation in varying speech styles and emotional tones. We analyzed user engagement patterns, perceived usability, emotional responses, and learning behaviors, with particular attention to how agent stylization influenced interaction across language proficiency levels and cultural backgrounds. Our findings reveal that agent design, especially voice, persona, and linguistic style, substantially affected user experience, motivation, and strategy. This work contributes to the understanding of affective, culturally stylized agents in human–agent interaction and offers guidance for designing more engaging, socially responsive systems.
Human-computer interaction, Computer-assisted language learning, Speech Processing
copyright: noneconference: ; ;
Figure 1. A screenshot from our Japanese language learning app, Jouzu, featuring “Hikari,” an anime-inspired conversational agent. Users interact via text (”send”) or voice (”record”), receiving stylized, in-character responses designed to enhance engagement and cultural immersion.
[…]
Screenshot of a conversation between the user and an anime-style agent named Hikari. The user asks for a Japanese word, and the agent replies in Japanese with a stylized, character-specific response.
1. Introduction
As conversational agents become integrated into everyday tools, the stylization of an agent’s voice, character, and aesthetics have become essential to the user experience. In human–agent interaction (HAI), conversational agents are not just tools, but social actors. They express personality, mediate emotion, and structure interaction in ways that influence user trust, engagement, and learning. This paper examines how stylistic and affective features of Japanese learning agents, particularly conversational characters, impact user interaction with the agents and user’s Japanese learning outcomes.
We focus on a multimodal interface where users engage with fictional agents designed to emulate recognizable character archetypes from anime, which is a form of Japanese animation that has recently become ingrained in pop culture around the world. By leveraging large language models (LLMs) and an expressive multilingual text-to-speech (TTS) system using Style-BERT-VITS2 (Rackauckas and Hirschberg, 2025), we produce emotionally rich, character-specific Japanese text and speech responses in conversational character agents. Unlike neutral chatbots, these agents employ language with non-standard pronouns, stylistic particles, and expressive phrasing to simulate social presence and emotional realism (Kinsui and Yamakido, 2015). Through these features, the agents act as affective, narrative participants who provide an amusing and engaging conversational environment rather than a traditional, robotic chatbot tutor.
Jouzu [^1] https://jouzu.ai features not only conversational characters, but also prescripted question-answer (QA) dialogues, or lessons, an in-text sentence-level inspector that defines and provides the readings of Japanese words, and flashcards to practice words learned in the lessons and conversational environments.
In a mixed-methods study with 54 participants, we evaluate how users of varying proficiency and linguistic backgrounds interacted with this application. We collected quantitative ratings of usability, engagement, and agent perception, alongside open-ended feedback and behavioral data from chatbot logs. In this study, we analyze how agent stylization such as voice, visual design, and character behavior modulates user interaction, emotional response, and task engagement. This contributes to ongoing discourse in HAI on the role of personality, culture, and embodiment in agent design.
2. Referenced Works
Language learning systems increasingly incorporate conversational agents to simulate spoken practice in a low-risk environment. Traditional computer-assisted language learning (CALL) tools have relied on scripted responses and task-based exercises, but newer systems use LLMs to support open-ended dialogue, generate feedback, and scaffold instruction (Cao and Zhong, 2023; Dornburg and Davin, 2024). Question-answer chatbots have been deployed in multiple languages to provide real-time practice, although most remain text-based (Li et al., 2022). More recent work has incorporated empathetic language to reduce anxiety and increase user retention (Siyan et al., 2024a, b). Aiba et al. (Aiba et al., 2024) introduced a system in which English learners could practice conference-style questions with GPT-generated speech, suggesting that combining TTS with LLMs encourages natural spoken interaction.
Voiced characters extend this trend by merging LLM dialogue with expressive speech synthesis and cultural stylization. While most prior systems employ neutral or utilitarian speech, Jouzu introduces character-specific language rooted in anime tropes—offering linguistic diversity and affective realism. These characters use “character language” featuring stylized pronouns, emphatic particles, and emotionally expressive phrasing, contributing to both social presence and learner motivation.
Kinsui and Yamakido (Kinsui and Yamakido, 2015) describe “character language” as a stylized form of Japanese used in fiction to signal personality, gender, or origin. These features include non-standard pronouns, archaic verb forms, and distinctive particles. Although rare in textbooks, they are recognizable and emotionally charged for many learners. Li et al. (Li et al., 2023) explored this concept with ChatHaruhi, a chatbot modeled on anime characters, but did not include voice interaction.
In contrast, Jouzu introduces voiced agents with character-specific linguistic stylization rooted in anime tropes. These agents use “character language” (Kinsui and Yamakido, 2015) to emulate fictional personas with affective realism. This represents a shift from purely pedagogical interactions toward immersive, socially engaging dialogue that blends linguistic and cultural learning.
Multimodal interfaces are central to current human-computer interaction (HCI) discussions about language learning. OpenAI’s GPT-4o (et al., 2024), for example, enables faster response generation with integrated audio (Ying et al., 2024), supporting smooth, responsive dialogue. At the same time, systems must address concerns about equity, safety, and engagement. Chan and Tsi (Chan and Tsi, 2023) note that computer-assisted tutors must maintain educational integrity while sustaining user interest. Xue et al. (Xue et al., 2023) highlight the importance of filtering harmful outputs in generative systems.
Previous HCI research has leveraged user studies to investigate and proposed guidelines for interactions with learner dialogue systems. For example, through a user study of 107 users, Koyoturk et al. assert the need for effective prompting in learner chatbot interactions, highlighting the need to provide user guidelines to learn how to prompt the agents (Koyuturk et al., 2025). Cai et al. conducted a user study with a group of three undergraduate students to advance dialogue system understanding in small group learning (Cai et al., 2024). Annamalai et al. conducted a user study with 360 undergraduate students to investigate chatbot language learning in Malaysian higher education institutions (Annamalai et al., 2023). Finally, Li et al. conducted a user study with 24 students to assess a curriculum-driven chatbot to match students’ classroom curricula (Li et al., 2024). Similarly, we employ a user study with 52 users to analyze user interactions and learning outcomes after their use of Jouzu, centered around the conversational character agents.
3. Method
A user study was conducted with 54 participants initially recruited; of these, 52 fully completed all parts of the study, including surveys and app usage. The analysis below focuses on these 52 complete cases to ensure consistency and validity. We examine app usability, learning outcomes, engagement, impact of stylized characters on the learning experience, and to compare Jouzu to other language learning apps like Duolingo. In addition, we gathered qualitative feedback on user experiences and analyzed interaction patterns between users and conversational agents.
Participants were recruited from a university computer science class. They were then instructed to freely use the app for at least one hour over a period of two weeks. In addition, we asked them to use the conversational agents present in the app. Users completed a survey consisting of multiple choice questions and open responses before (see Appendix A) and after the app (see Appendix B). The initial survey collected demographic and background information from users and presented an optional 10-question multiple choice Japanese quiz with an optional short paragraph response in Japanese (see Appendix C). The post-use survey asked several questions about the user experience, followed by a new 10-question Japanese quiz of the same difficulty level and an open response in Japanese (see Appendix C). Participants were rewarded with 5 assignment points in their university class. Of the 54 participants, 52 fully completed the task.
3.1. Descriptive Statistics of User Experience
For quantitative analysis, Likert scale questions were grouped into important aspects for evaluation of the app. The following scale was used for groups one through 5:
  1. (1)
    1: Not at all
  2. (2)
    2: Slightly
  3. (3)
    3: Somewhat
  4. (4)
    4: Moderately
  5. (5)
    5: Extremely
For group six:
  1. (1)
    1: Much worse
  2. (2)
    2: Slightly worse
  3. (3)
    3: The same
  4. (4)
    4: Slightly better
  5. (5)
    5: Much better
  1. (1)
    Usability
    • How easy was the app to use?
  2. (2)
    Amount Learned
    • How much did the app help you learn or practice Japanese?
  3. (3)
    Fun and Engagement
    • How engaging was the app overall?
    • How fun was it to use?
    • How motivated did you feel to keep using it?
    • How much did the animations enhance your engagement?
  4. (4)
    Characters
    • How much did the characters enhance your learning experience?
    • Which character did you enjoy chatting with the most?
    • Did the characters make the app more enjoyable overall?
  5. (5)
    Visuals, Art, and Aesthetic
    • How much did the anime-style visuals make the app more appealing?
  6. (6)
    Comparison to Other Apps
    • Compared to other language learning apps, how would you rate the app in terms of fun?
    • Compared to other language learning apps, how would you rate the app in terms of helpfulness for learning?
    • Compared to other language learning apps, how would you rate the app in terms of ease of use?
    • Compared to other language learning apps, how would you rate the app in terms of your motivation to keep learning?
    • Compared to other language learning apps, how would you rate the app in terms of visual and auditory style?
3.2. Hypotheses and Analysis Design
We tested two hypotheses for significance:
  1. (1)
    Native Chinese (Mandarin) users will learn more than users of other native backgrounds since Japanese and Chinese share Kanji characters.
  2. (2)
    Users with a higher level of Japanese will learn more and find the app more fun and engaging compared to beginners of Japanese due to Jouzu’s immersive approach.
To examine potential cross-linguistic differences in user experience, participants were grouped by native language: Chinese and non-Chinese speakers.
To assess the significance of these cross-linguistic differences, we conducted independent-sample t-tests comparing mean group ratings between Chinese and non-Chinese participants across all six experience categories.
We also grouped users by self-reported Japanese proficiency (Complete Beginner, Beginner, Intermediate, Native) to analyze mean ratings by language experience level. A heatmap was generated to visualize rating trends across these categories.
3.3. Qualitative Data Collection
Participants responded to four open-ended prompts:
  • What was the most helpful feature for learning Japanese?
  • What was the least helpful feature?
  • What was the most fun and engaging feature?
  • What did you find frustrating or confusing?
Responses were mapped to one of six predefined experience categories: Usability, Amount Learned, Fun and Engagement, Characters, Visuals, Art and Aesthetic, and Comparison to Other Apps.
3.4. Realized Learning Outcomes
3.4.1. Multiple Choice
Users were asked ten multiple choice questions before and after use of the app. All questions were optional. The questions are included in . Pre- and post-use completion rates and accuracy scores were compared to assess realized learning outcomes in terms of grammar and vocabulary.
Ten questions were asked. The first five questions included vocabulary sampled from the characters’ individual language styles. The next five were questions sampled from each Japanese Language Proficiency Test Level (N1 through N5).
The pre- and post-tests were different, but were created to be the same difficulty for respondents.
3.4.2. Open Paragraph
Participants were invited to write a short paragraph in Japanese both before and after using the app. This task was optional and open-ended, allowing users at any level to attempt expressive written output. Pre- and post-use submissions were compared for fluency, vocabulary diversity, and structural complexity.
3.5. Chatbot Interaction Logging
User interaction data with the character-based chatbot was analyzed qualitatively. Patterns were identified in beginner vs. advanced usage, including:
  • English queries for vocabulary (e.g., “What is ‘sword’ in Japanese?”)
  • Passive listening via English prompts with Japanese replies
  • Full-sentence Japanese dialogue from intermediate and native speakers
  • Use of speech input for pronunciation and listening practice
These patterns were coded from logged conversations or retrospective user descriptions.
4. Analysis
4.1. User Background
In the pre-use survey, background information including native language, other languages spoken, age, and other language learning apps used was collected from users. Below, we present the background information of the users who completed the study and provided background information: 50 participants.
English was the most frequently reported native language, accounting for 19 responses, with bilingual variants such as Korean and English present. Other native languages included Chinese (17), Korean (3), Hindi (5), Tamil (1), urkish (1), Japanese (1), and Malayam (1), reflecting a linguistically diverse sample. The age distribution showed a concentration of younger participants, with an average age of 24.32 years (SD = 3.78). The age range extended from 20 to 43 years, with 50% of respondents aged 23 or younger and 75% under the age of 25, indicating that most participants were in their early twenties. 39 participants had previously used Duolingo. Other language learning apps used included Rosetta Stone (4), Busuu (3), Babbel (3), and Memrise (1).
4.2. Qualitative Analysis
Participant responses were aggregated across six experience dimensions. Usability received the highest average rating (M = 3.64), followed closely by Visuals, Art, and Aesthetic (M = 3.52) and Characters (M = 3.39), indicating that users responded positively to the interface design and stylized agents. Fun and Engagement also scored well overall (M = 3.31, SD = 0.36), though variation in motivation and enjoyment was observed. Comparison to other apps showed moderate ratings (M = 3.10, SD = 0.46), reflecting mixed perceptions of how Jouzu compares to traditional language learning tools. Perceived learning outcomes were lowest (M = 2.48), suggesting that while the app was enjoyable and visually engaging, users may not have viewed it as highly effective for structured language acquisition.
Figure 2 shows the average rating per group with standard deviation error bars, highlighting the overall trends and relative spread between each category. Table 1 presents the full descriptive statistics by group.
Figure 2. Mean participant ratings by experience category with standard deviation error bars (Likert scale: 1–5).
Table 1. Group-level descriptive statistics of user ratings (Likert scale: 1–5)
Group Mean SD Min Max Avg Item SD N
Amount Learned 2.48 2.48 2.48 1.09 44
Characters 3.39 3.39 3.39 1.48 44
Comparison to Other Apps 3.10 0.46 2.51 3.74 1.12 214
Fun and Engagement 3.31 0.36 2.77 3.55 1.13 176
Usability 3.64 3.64 3.64 1.10 44
Visuals, Art, and Aesthetic 3.52 3.52 3.52 1.37 44
4.3. Demographic Significance
We tested two hypothesis for significance:
  1. (1)
    Native Mandarin users will learn more than users of other native backgrounds since Japanese and Chinese share Kanji characters.
  2. (2)
    Users with a higher level of Japanese will learn more and find the app more fun and engaging compared to beginners of Japanese due to Jouzu’s immersive approach.
4.3.1. Native Mandarin User Ratings
To examine potential cross-linguistic differences in user experience, participants were grouped by native language: Chinese (including Mandarin) and non-Chinese speakers. Across all six dimensions, Chinese speakers rated the app more favorably. Most notably, the Characters group showed a large gap (M = 4.29 vs. 2.97), suggesting strong engagement with stylized agents among Chinese users. Visuals, Art, and Aesthetic also received a much higher rating from Chinese speakers (M = 4.29 vs. 3.17), followed by higher scores for Fun and Engagement (M = 3.82 vs. 3.07). Usability and Comparison to Other Apps showed smaller but consistent differences, while the largest perceived gap in learning outcomes (M = 2.79 vs. 2.33) aligned with open-ended feedback. These results point to meaningful cultural or pedagogical variation in how users interact with character-driven language tools. See Table 2.
Figure 3 visualizes the mean rating differences between Chinese and non-Chinese participants across all experience groups, highlighting consistently higher ratings by Chinese users.
Figure 3. Mean group ratings for Chinese vs. non-Chinese native speakers.
Table 2. Mean experience ratings by native language group
Group Chinese Mean Non-Chinese Mean
Amount Learned 2.79 2.33
Characters 4.29 2.97
Comparison to Other Apps 3.45 2.95
Fun and Engagement 3.82 3.07
Usability 3.86 3.53
Visuals, Art, and Aesthetic 4.29 3.17
To assess the significance of these cross-linguistic differences, we conducted independent-sample t-tests comparing mean group ratings between Chinese and non-Chinese participants across all six experience categories. As shown in Table 3, Chinese participants gave significantly higher ratings in the areas of Characters, Fun and Engagement, Visuals, and Comparison to Other Apps. These results suggest that the stylized elements of Jouzu, particularly character-driven dialogue and anime aesthetics, may resonate more strongly with learners from Chinese-language backgrounds. While differences in Usability and Amount Learned did not reach statistical significance, both trended in the same direction, indicating a broader pattern of increased satisfaction among Chinese-speaking users.
Based on this significance testing, our first hypothesis was not supported. However, native Chinese speakers performed significantly better for all groups in addition to the amount of learned and usability. That is, native Chinese speakers found the app more fun and engaging, liked the inclusion of anime characters better, preferred the visuals, art, and aesthetic, and found it better overall compared to other language learning apps.
Table 3. Independent-sample t-tests by native language group
Group CN Mean NCN Mean t p NCN NNCN
Amount Learned 2.79 2.33 1.49 .1459 14 30
Characters 4.29 2.97 3.27 .0026 14 30
Comparison to Other Apps 3.45 2.95 2.92 .0042 65 149
Fun and Engagement 3.82 3.07 4.48 <<<.0001 56 120
Usability 3.86 3.53 0.98 .3359 14 30
Visuals, Art, and Aesthetic 4.29 3.17 2.95 .0059 14 30
4.3.2. Differences by Japanese Proficiency Level
To better understand how language experience shapes user impressions, we analyzed mean group ratings across self-reported Japanese proficiency levels, ranging from complete beginners to near-native speakers. As shown in Table 4, user satisfaction generally increased with proficiency. Intermediate and native-level users rated Characters and Fun and Engagement highest, both scoring above 3.8 on average. In contrast, complete beginners rated these same categories lower (3.12 and 3.24, respectively), suggesting that more fluent users were better able to engage with stylized dialogue and game-play elements.
A notable difference was observed in the Characters category, where Beginner users reported a sharp increase in enjoyment (M = 4.25) compared to Complete Beginners (M = 3.12), possibly reflecting increased comprehension of nuanced dialogue. Visuals, Art, and Aesthetic also saw higher scores from more advanced users, echoing trends seen in Figure 2. Overall, these results indicate that Jouzu may provide greater value and enjoyment as learners advance in proficiency.
Table 4. Mean experience ratings by Japanese proficiency level
Group Complete Beginner Beginner Intermediate Native
Amount Learned 2.39 2.62 3.00 3.00
Characters 3.12 4.25 4.00 4.00
Comparison to Other Apps 2.96 3.42 3.80 3.80
Fun and Engagement 3.24 3.41 3.88 3.88
Usability 3.70 3.12 4.50 4.50
Visuals, Art, and Aesthetic 3.39 3.88 4.00 4.00
Figure 4. Mean group ratings by Japanese proficiency level.
A heatmap of mean ratings across Japanese proficiency levels (Figure 5) reveals clear trends in how learners with different backgrounds experienced the app. Ratings for all categories increased with proficiency, but the most pronounced differences appeared in Characters, Fun and Engagement, and Visuals, Art, and Aesthetic. Complete beginners rated the character experience at M = 3.12, while intermediate and native-level users both gave it M = 4.00 or higher. Similar patterns held for engagement and visual appeal, suggesting that stylized, character-driven interactions resonated more strongly with users who could understand more of the dialogue and cultural nuance. In contrast, usability showed less consistent variation, indicating that learners at all levels found the app interface accessible. These trends underscore the importance of matching language level to interaction design in character-based learning environments.
Figure 5. Heatmap of mean experience ratings across Japanese proficiency levels. Darker shades indicate higher user satisfaction per category.
5. Qualitative Feedback
5.1. Qualitative Feedback by Experience Group
We conducted a qualitative analysis of open-ended survey responses to four short-answer prompts, using a structured keyword mapping to associate each response with one of six predefined experience categories: Usability, Amount Learned, Fun and Engagement, Characters, Visuals, Art, and Aesthetic, and Comparison to Other Apps. The four prompts asked participants to describe (1) the most helpful feature for learning Japanese, (2) the least helpful feature, (3) the most fun and engaging feature, and (4) any aspects they found frustrating or confusing.
For the question about the most helpful feature, the majority of responses were associated with the Amount Learned group. Participants frequently praised the chatbot and quiz mechanics for their ability to reinforce vocabulary and simulate real conversations. One user explained, “The chat feature was honestly the most helpful because it essentially was a chatbot companion for my learning,” while another appreciated being able to ask for phrases and practice pronunciation interactively.
When asked about the least helpful feature, responses were again concentrated in the Amount Learned group, but with a more critical tone. Users cited difficulty understanding advanced responses in the chat, a lack of scaffolding for beginners, and confusion about how to progress. One participant noted, “You need to have a basis of how to read/write and understand Japanese. Something more simple that starts with the Japanese alphabet would be more helpful.” Usability concerns also surfaced, including confusion in the Learn feature and overwhelming text density in early lessons.
Responses to the most fun and engaging feature were overwhelmingly mapped to the Fun and Engagement category. Learners described enjoyment from story interactions, chatting with voiced characters, and the gamified nature of lessons. One user remarked, “It felt like chatting with a friendly sensei (teacher)… It wasn’t just study, it felt like real communication.” Others highlighted specific character personalities and the immersive experience of interacting with anime-like avatars.
Finally, in response to frustrations and confusion, feedback was mostly provided for Usability and Characters. Users reported UI bugs, confusing navigation, lack of direction, and speech recognition issues. One respondent wrote, “I didn’t know how to use the flash cards and couldn’t figure it out,” while another explained that character voices were too quiet and hard to understand. A few users also criticized the app’s difficulty curve, citing challenges understanding kanji and parsing lesson content without sufficient guidance.
Overall, these qualitative results reinforce the quantitative trends observed earlier: participants with more language experience tended to appreciate the app’s depth and character-driven content, while beginners expressed a need for greater structure, onboarding, and transparency in lesson progression.
5.2. Realized Change in Japanese Proficiency
5.2.1. Multiple Choice Performance Analysis
This analysis compares participant performance across two sets of graded Japanese language questions before and after app usage. Among participants who completed both quizzes, average scores decreased slightly from 6.50 to 4.76 out of 10. However, high-end performance remained stable: two participants achieved perfect scores in the post-test, and six scored eight or higher. This suggests that, while some users may have struggled with retention or quiz difficulty, others maintained or improved their performance. These results, combined with increased comfort shown in open-ended output, indicate that the app may have supported language exploration and engagement, even if structured gains varied.
Metric (Attempting Users Only) First Set Second Set
Participants Responded 16 34
Average Score 6.50 4.76
Perfect Scores (10/10) 0 (0.0%) 2 (5.9%)
High Scores (8–10) 2 (12.5%) 6 (17.6%)
Lower Scores (1–3) 5 (31.3%) 14 (41.2%)
Table 5. Comparison of engagement and performance before and after app use (among attempting users)
5.2.2. Open Paragraph
Users were instructed to optionally write a paragraph in Japanese about anything they wished to the best of their ability. Before using the app, five respondents of varying proficiency levels – beginner, intermediate, and native – wrote paragraphs in Japanese. For example, an intermediate speaker wrote the following:
{CJK}
UTF8min
私は日本語を勉強しています. 最初はひらがなとカタカナを覚えるのが難しかったです.でも, 毎日少しずつ練習して, 今は簡単な会話ができます. 将来,日本に旅行して, 実際に日本語を使ってみたいです.
Translation: I am studying Japanese. At first, it was difficult to memorize Hiragana and Katakana. But by practicing little by little every day, I can now have simple conversations. In the future, I want to travel to Japan and try using Japanese in real life.
In contrast, after usage of the app, nine students wrote a paragraph in Japanese. Of these nine students, three had previously self-identified as complete beginners with no exposure to Japanese.
The paragraphs from participants who were unable to write them before had varying degrees of fluency. For example, one student who identified as a beginner wrote the following:
{CJK}
UTF8min
私の日本語はまだ初歩的ですが、食べ物などについて書けるようになりました。うどんは歯ごたえがあってスープが美味しいので大好きです。近いうちに日本に行くのが楽しみです。大学卒業後かな。
Translation: My Japanese is still basic, but I’ve become able to write about things like food. I love udon because it has a nice chewy texture and the soup is delicious. I’m looking forward to going to Japan sometime soon—maybe after I graduate from university.
Others demonstrated basic vocabulary acquisition, not limited to the vocabulary words used in the character’s individual language styles.
{CJK}
UTF8min
konichiwa, ogenki desu ka. Sumimasen, doko toire desu ka. Sore nan desu ka. Arigatou. Naze (why), itsu (when), dare (who).
Translation: Hello, how are you? Excuse me, where is the bathroom? What is that? Thank you. Why, when, who.
Another student:
おはようございます- お元気ですか (My attempt at ”good morning, how are you?”)
One participant who was already identified as a beginner with some experience did not attempt a paragraph before use, but constructed a comprehensive sentence after use. It is important to note that, while characters do not use polite Japanese, this student still wrote a sentence using polite Japanese mixed with Saya’s character language.
{CJK}
UTF8min
私は武道が好きです
Translation: I like bushi
5.3. Chatbot Interaction
We observed diverse interaction strategies among users engaging with Jouzu’s character-based chat feature. Across early-stage usage, three recurring patterns emerged, particularly among beginners.
First, several participants initiated conversations in English by asking the characters to introduce a Japanese word. In response, characters typically presented a common word or phrase, accompanied by a simple explanation in Japanese. For example, one user typed, “Teach me a word in Japanese,” prompting a character to reply with an entry-level greeting and its meaning. These interactions provided accessible entry points into vocabulary acquisition, even for those with no prior exposure to Japanese.
Second, users frequently asked how to translate specific English words into Japanese, such as “How do you say sword in Japanese?” Character responses often included the target term in Kanji with furigana or spoken pronunciation, demonstrating how the system scaffolds lexical learning through bilingual dialogue. This behavior echoes documented beginner strategies in computer-assisted language learning, where learners rely on translation-based queries to build foundational vocabulary.
Third, a subset of participants conducted simple, English-initiated conversations about everyday topics. The characters responded entirely in Japanese, allowing users to passively absorb sentence structure and phrasing. For instance, a user might ask, “What is your favorite tea?” and receive a culturally contextualized response in natural-sounding Japanese. These exchanges continued smoothly, with characters maintaining conversational coherence even when users offered minimal or English-only input.
In contrast, intermediate and advanced users tended to rely more heavily on Japanese in their chat entries. Many incorporated full sentences and questions in Japanese, using the characters to simulate real-time dialogue in a stylized but grammatically accurate register. Notably, more fluent users also engaged with the speech recognition functionality, using spoken input to practice pronunciation and listening comprehension. These behaviors suggest that the chat system offers scalable interaction, adapting fluidly to user proficiency levels. While beginners benefit from bilingual support and forgiving prompts, advanced users appear to leverage the system as a semi-authentic environment for spoken Japanese practice.
6. Results
We conducted a mixed-methods evaluation of user interaction with Jouzu, a multimodal language learning environment featuring stylized, voiced conversational character agents. Participants rated the system most highly for usability (M = 3.64), visual and aesthetic appeal (M = 3.52), and character engagement (M = 3.39), indicating strong positive appreciation of both the interface design and affectively rich agent behavior. Although perceived learning outcomes received lower scores on average (M = 2.48), open-ended responses suggest that users viewed Jouzu primarily as a motivational, immersive supplement to traditional instruction, rather than a standalone pedagogical solution.
Demographic analysis revealed consistent trends. Native Chinese speakers rated the system significantly higher across multiple experience dimensions, particularly for visual appeal, character interaction, and overall engagement. This pattern suggests stronger alignment with the anime-inspired aesthetic and the affective realism of stylized agent dialogue. Additionally, participants with higher self-reported Japanese proficiency gave more favorable ratings across all categories. Intermediate and native-level users in particular reported greater enjoyment of character language, voiced responses, and narrative immersion, reinforcing the utility of stylization at more advanced stages of language learning.
Analysis of interaction logs further revealed differentiated usage strategies by proficiency level. Complete beginners commonly used English prompts to request translations or vocabulary (e.g., “What is ‘sword’ in Japanese?”), with agents responding in Japanese using character-specific phrasing. These users often relied on passive listening or hybrid input patterns. In contrast, intermediate and advanced users engaged in full-sentence Japanese conversations, using both text and speech input to simulate native dialogue interaction.
Writing samples submitted before and after app usage revealed meaningful expressive gains, particularly among novice learners. Several users who were unable to produce any Japanese output prior to using the app submitted short paragraphs afterwards, indicating increased confidence and productive vocabulary. These paragraphs included both basic grammatical constructions and character language vocabulary, suggesting exposure to stylistic features embedded in the agent dialogue.
Taken together, these findings demonstrate that Jouzu’s stylized conversational agents foster high levels of engagement and perceived usability across a diverse user base. While advanced learners derived the most direct value from immersive dialogue, even complete beginners demonstrated affective benefits and signs of early language uptake.
7. Discussion
Our findings suggest that stylized, voiced dialogue agents can significantly enhance user engagement and motivation in computer-assisted language learning. While Jouzu offers multiple learning features including lessons, flashcards, and a sentence-level inspector, the emotionally expressive agents were consistently cited as the most engaging component. Users reported that the character voices, visual design, and in-character phrasing created a more immersive and socially meaningful learning experience than traditional tools.
However, stylization alone was not sufficient for all users. Beginners often struggled to understand the agents’ responses without more structured support, highlighting the need for adaptive scaffolding and clearer progression cues. In contrast, intermediate and advanced learners reported higher enjoyment and made fuller use of the conversational agents, sometimes mirroring their stylized language. These patterns reflect existing work on personalization in dialogue and CALL systems and show that character-based interaction may be most effective when aligned with the learner’s proficiency.
While multiple choice test scores decreased slightly after app usage, this does not necessarily indicate a decline in learning outcomes. Jouzu emphasizes open-ended, affective interaction over rote memorization, and several users who showed weaker multiple choice performance nonetheless demonstrated stronger expressive ability in paragraph writing. Many participants who initially wrote little or no Japanese were later able to produce multi-sentence output using vocabulary, grammar, or stylistic features encountered in the app. This suggests that the app supports exploratory, confidence-building learning, which may not be fully captured by short-format assessments.
Overall, Jouzu’s design suggests that culturally stylized, emotionally expressive dialogue agents—when embedded in a multimodal learning environment—can drive engagement and support meaningful language practice. Future work should explore adaptive features to help beginner users benefit more fully from the system’s immersive elements.
8. Conclusion
Jouzu demonstrates that stylized, character-based interaction can enhance user engagement and cultural immersion in language learning, though structured learning support remains essential for beginner users. Future work will explore adaptive scaffolding for novice learners and investigate long-term impacts of character-driven dialogue on spoken fluency.
Acknowledgements
This research was initiated independently of the primary author’s academic responsibilities at Columbia University and was conducted without the use of university resources.
{CJK}
UTF8min
Appendix A Survey Questions
A.1. Pre-use Questions
The users were introduced to the survey with the following introduction.
Welcome to my study! This study is a user study for Jouzu, a mobile application for learning Japanese with original anime-inspired characters. This survey will ask for basic information regarding your device, and a 10-15 minute assessment in Japanese. Your UNI (for CU students) and/or email (if external) will be tracked for assignment credit purposes. All data will be anonymized for analysis and publication. Please note that refusal to participate will have no influence on your academic standing.
By signing up for this user test and creating an account on the app, you agree to our Terms of Service.
The users were asked the following background information questions.
  1. (1)
    Have you studied Japanese before? (It’s completely okay if you haven’t)
  2. (2)
    What is your native language?
  3. (3)
    Which other languages do you speak? (You don’t have to be completely fluent in them, but at least have a basic understanding).
  4. (4)
    What is your age?
  5. (5)
    Which (if any) of these apps have you used before?
  6. (6)
    Which (if any) of these games have you played?
  7. (7)
    Which (if any) of the following games have you played?
Appendix B Appendix: App Evaluation Questions
Users were introduced to the post-survey with the following introduction.
Thank you very much for using my app! Here, you will be asked several questions about your experience, then take another short Japanese assessment. Please answer all questions to the best of your ability this time, but don’t panic if you don’t know any answers. There is no expectation for your score.
The following questions were asked to evaluate the application.
  1. (1)
    How long did you use the app?
  2. (2)
    Which features did you use?
  3. (3)
    How engaging was the app overall?
  4. (4)
    How fun was it to use?
  5. (5)
    How much did the app help you learn or practice Japanese?
  6. (6)
    How motivated did you feel to keep using it?
  7. (7)
    How easy was the app to use?
  8. (8)
    Which feature was the most helpful for you to learn and practice Japanese?
  9. (9)
    Which feature was the least helpful for you to learn and practice Japanese?
  10. (10)
    Which feature was the most frustrating or confusing?
  11. (11)
    Which feature was the most fun and engaging?
  12. (12)
    How much did the music improve your experience?
  13. (13)
    How much did the animations enhance your engagement?
  14. (14)
    How much did the anime-style visuals make the app more appealing?
  15. (15)
    How much did the characters enhance your learning experience?
  16. (16)
    Which of these aesthetic elements stood out the most? (Short answer)
  17. (17)
    Which character did you enjoy chatting with the most?
  18. (18)
    Why did you prefer that character? (Short answer)
  19. (19)
    Did the characters make the app more enjoyable overall?
    1 = not at all, 2 = slightly, 3 = somewhat, 4 = moderately, 5 = extremely
  20. (20)
    Compared to other language learning apps, how would you rate the app in terms of fun?
    1 = much worse, 2 = slightly worse, 3 = the same, 4 = slightly better, 5 = much better
  21. (21)
    Compared to other language learning apps, how would you rate the app in terms of helpfulness for learning?
    1 = much worse, 2 = slightly worse, 3 = the same, 4 = slightly better, 5 = much better
  22. (22)
    Compared to other language learning apps, how would you rate the app in terms of ease of use?
    1 = much worse, 2 = slightly worse, 3 = the same, 4 = slightly better, 5 = much better
  23. (23)
    Compared to other language learning apps, how would you rate the app in terms of your motivation to keep learning?
    1 = much worse, 2 = slightly worse, 3 = the same, 4 = slightly better, 5 = much better
  24. (24)
    Compared to other language learning apps, how would you rate the app in terms of visual and auditory style?
    1 = much worse, 2 = slightly worse, 3 = the same, 4 = slightly better, 5 = much better
  25. (25)
    What did you like most about the app?
  26. (26)
    What did you like least or find frustrating/
    confusing?
  27. (27)
    What would you change about the app?
  28. (28)
    Any suggestions for improvement?
Appendix C Pre- and post-use Japanese Quiz Questions
C.1. Pre-use Quiz
The users were introduced to the Japanese quiz with the following explanation.
Here, you will be asked a few questions to assess your Japanese knowledge. The results of this test will be compared against another test after you finish using the app.
If you have not studied Japanese before or are not confident in taking a Japanese test, please skip this part and submit the form. Skipping this part will not affect your participation or assignment credit in any way.
For those who decide to continue, if you don’t know an answer, feel free to skip the question or take your best guess. Some of these questions will be very difficult and may be beyond your current knowledge, so do not stress about getting a high score.
The users were asked the following questions.
  1. (1)
    「しんしん」を漢字で書いてください。
  2. (2)
    紡ぎ(つむぎ)の意味を選びなさい。
  3. (3)
    「じひ」を漢字で書いてください。
  4. (4)
    友達(ともだち)とは何ですか?
  5. (5)
    語らい(かたらい)とは何ですか?
  6. (6)
    新しいくるまですね
  7. (7)
    (  )しないで、どうぞたくさんたべてください。
  8. (8)
    3日前(みっかまえ)から雨(あめ)が続いている。
  9. (9)
    事故の原因は、機械の(  )作動にあると考えられている。
  10. (10)
    自分が未熟すぎて心配だなどというのは、未熟ということをマイナスに考える証拠だ。(  )弱い人間とか未熟な人間のほうが、はるかにふくれあがる可能性をもっている。
  11. (11)
    Please write 1 paragraph (at least 4 sentences) in Japanese about any topic you would like.
C.2. Post-use Quiz
The users were introduced to the post-use quiz with the following instructions.
Please answer as many questions as you can to the best of your ability. None are required, however, so do not feel pressured to answer them if you do not know the answers.
The following quiz questions were asked to the user.
  1. (1)
    精神(せいしん)とは何ですか?
  2. (2)
    名誉を読むとき、正しいふりがなは何ですか?
  3. (3)
    崇高(すうこう)に最も近い英語の意味を選んでください。
  4. (4)
    自然観を読むとき、正しいふりがなは何ですか?
  5. (5)
    歴史(れきし)の意味を選んでください。
  6. (6)
    きのうはうちに(  )何をしましたか。
  7. (7)
    今日は、ゴミのしゅうしゅう日ですか。
  8. (8)
    わたしのせんもんは医学です。
  9. (9)
    登山(とざん)の途中(とちゅう)で、(  )と思いました(おもいました)。なぜかというと、予想(よそう)よりも山(やま)の道(みち)を歩く(あるく)のは大変(たいへん)だったからです。
  10. (10)
    おもちゃというと、ただ子供が遊ぶためだけのものだと(  )。
  11. (11)
    Please write 1 paragraph in Japanese about any topic you would like.